OpenStudy (jokersmart):

Which is equal to 6^-8/6^2? A. 616 B. 64 C. 6−6 <--? D. 6−10

OpenStudy (anonymous):

\[-8-2=-10\]

OpenStudy (anonymous):

that is a different story

OpenStudy (jokersmart):

What? o-o

OpenStudy (anonymous):

are both exponents positive, or it \[\huge \frac{6^{-8}}{6^2}\]

OpenStudy (jokersmart):

They're like this 

OpenStudy (jokersmart):

I think negative

OpenStudy (anonymous):

oh then your answer is right \[-8-(-2)=-6\]

OpenStudy (jokersmart):

Okay, thanks!